OpenStudy (anonymous):

write the multiplication and divsion facts for the fact family with the number 5,6,and 30

OpenStudy (anonymous):

5*6=30 and 6*5=30 and 30/5=6 and 30/6=5....... Is that what you meant?

OpenStudy (anonymous):

i think so its 4th grade work lol iam helping my cuzo